Which of the ordered pairs
(3, 1), (0, –4), (–4, 0), (–3, –7)
are solutions for the equation x – y = 4?
A) (0, –4), (–4, 0), and (–3, –7)
B) (–4, 0) and (–3, –7)
C) (0, –4) and (–3, –7)
D) (3, 1) and (–4, 0)
No book. Thank you so very much for your time!!!

Answer by rchill(405)

For each ordered pair, substitute the values of x and y into the equation to see if the statement is true. If the statement is true, then the ordered pair is a solution.

(3,1): 3-1=4, which becomes 2=4, which is FALSE
(0,-4): 0-(-4)=4, which becomes 4=4, which is TRUE
(-4,0): -4-0=4, which becomes -4=4, which is FALSE
(-3,-7): -3+7=4, which becomes 4=4, which is TRUE

So, the ordered pairs that are solutions are (0,-4) and (-3,-7), or answer (C).